COMMISSIONING ACTIVITIES FOR THE ACE MISSION

\underline{RAY NASSAR}, CHRIS BOONE, KALEY A. WALKER, SEAN D. MCLEOD, PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada, N2L 3G1.

The SCISAT-1 satellite for the Atmospheric Chemistry Experiment (ACE) mission was launched on August 12, 2003. The main goal of the mission is to investigate the chemical and dynamical processes that control the distribution of ozone in the stratosphere and upper troposphere. The ACE instruments record spectra of Earth’s atmosphere primarily using the solar occultation method. Commissioning activities and on-orbit data from the high resolution infrared Fourier transform spectrometer (ACE-FTS), the UV-visible-NIR MAESTRO spectrometer and the ACE solar imagers will be explained with emphasis on the ACE-FTS.